Mary is 5 times as old as Jamie. in 10 years Mary will be three times as old as Jamie, how old is each now?

Answer by jorel1380(3719)

Let j be Jamie. Then Mary would be 5j. So:
5j+10=3(j+10)
5j+10=3j+30
2j=20
j=10
Jamie is 10; Mary is 50. ☺☺☺☺